OpenStudy (anonymous):

7 3/5 + 3 3/4 answer choices: a. 11 1/3 b. 11 7/20 c. 7 1/3 d.10 1/3

OpenStudy (cwrw238):

3/5 + 3/4 lcm = 20: 12/20 + 15/20 = 27/20 = 1 7/20 7 + 3 + 1 7/20 = 11 7/20

OpenStudy (amorfide):

just to help incase you don't understand how to add fractions lets say we have 2 fractions a/b + c/d